VIBRATIONAL PREDISSOCIATION DYNAMICS OF THE (H_{2}O)_{2} DIMER

\underline{L. C. CH’NG}, B. E. ROCHER, A. K. MOLLNER, and H. REISLER, Department of Chemistry, University of Southern California, Los Angeles, CA, 90089.

The state-to-state vibrational predissociation dynamics of the (H_{2}O)_{2} dimer were studied by resonance-enhanced multiphoton ionization (REMPI) and velocity-map imaging (VMI) to obtain pair-correlated product energy distributions. The 2+1 REMPI spectra of the H_{2}O photofragments were recorded via the C̃^{1}B_{1} (000) ← X̃^{1}A_{1} (000 and 010) transition following a vibrational excitation of the dimers bound-OH stretch fundamental. The fragment center-of-mass translational energy (c.m. E_{T}) distributions were determined from VMI of selected rotational states of the detected H_{2}O photofragments. The c.m. E_{T} distributions were then converted to pair-correlated H_{2}O cofragment rotational level distributions. This is the first experiment in which H_{2}O products with bend (ν_{2}) excitation were observed by REMPI. The dissociation energy of the dimer was determined from the images with spectroscopic accuracy. The predissociation mechanism of (H_{2}O)_{2} will be discussed and compared with the corresponding hydrogen bonded dimers of an acid (HCl-H_{2}O) and a base (NH_{3}-H_{2}O).